OpenStudy (anonymous):

graph the line containing the given pair of point and find the slope (6,5),(-6,-3)

OpenStudy (anonymous):

\[\Delta y=5-(-3)=8\]\[\Delta x=6-(-6)=12\]Thus the slope is\[y=\Delta y/\Delta x=8/12=2/3\]

OpenStudy (anonymous):

Should be m in the last line (sorry)

OpenStudy (anonymous):

so m=8/12

OpenStudy (anonymous):

so how would i graph it ?

OpenStudy (anonymous):

The slope is Change in y / change in x So it's 2/3. The line equation would be y = 2/3x + 1

OpenStudy (anonymous):

There would be a point at (-3,-1), (0,1), and (3, 3).

OpenStudy (anonymous):

m=2/3